OpenStudy (anonymous):

What is the sum of a 12–term arithmetic sequence where the last term is 13 and the common difference is –10?

OpenStudy (amistre64):

if we knew a1 we would be set

OpenStudy (amistre64):

a12 = a1 +d(11) a12 - d(11) = a1

OpenStudy (amistre64):

\[S_{12}=\frac{n(a_1+a_{12})}{2}\] \[S_{12}=\frac{n(a_{12}-d(11)+a_{12})}{2}\] \[S_{12}=\frac{n(2a_{12}-d(11))}{2}\]